OpenStudy (anonymous):

What is the volume of a sphere with a surface area of 9pi yd^2? A) 9pi yd^3 B) 36pi yd^3 C) 9/2pi yd^3 D) 9/4pi yd^3

OpenStudy (amistre64):

surface area = 4pi r^2 right?

OpenStudy (amistre64):

volume = (4/3) pi r^3

OpenStudy (amistre64):

9 pi = 4 pi r^2 9 = 4 r^2 r = 3/2 then correct?

OpenStudy (amistre64):

\[Vol=\frac{4pi}{3}*(\frac{3}{2})^3\]

OpenStudy (amistre64):

that comes to 4.5 pi if i dint mess it up